What weight of lactic acid must be dissolved in a liter solution to produce a hydronium concentration of 1 x 10^-3 F?


1
Expert's answer
2021-03-22T06:12:31-0400

concentration or molarity = 1 ×10-3 M

Volume = 1L

Molarity = "\frac{no. Of moles of solute}{volume of solution}"


1×10-3 = "\frac{moles}{1}"


moles of lactic acid = 1×10-3 mol

weight of lactic acid = 1×10-3 ×90.08

= 90.08×10-3 = 0.09008 g